OpenStudy (anonymous):

log(5)70=____? round the final answer 4 decimal places.

OpenStudy (anonymous):

> log(5)*70 [1] 112.6607

OpenStudy (anonymous):

\[\text{Let } k = \log_5(70)\]\[\implies 5^k = 70\]\[\implies \log(5^k) = \log(70)\]\[\implies k\log(5) = \log(70)\]\[\implies k = \frac{\log(70)}{\log(5)}\]\[\implies \log_5(70) = \frac{\log(70)}{\log(5)}\]

OpenStudy (anonymous):

thanks, but that's incorrect apparently mapologo

OpenStudy (anonymous):

so....

OpenStudy (anonymous):

i just divide 70 by the base (5), and get 14... but what to do about the 'log' part??

OpenStudy (anonymous):

No. You divide the log of 70 by the log of the base.

OpenStudy (anonymous):

ugh...

OpenStudy (anonymous):

2.6397??

OpenStudy (anonymous):

It's the change of base 'formula' but I can't bother remembering it so I derive it each time.

OpenStudy (anonymous):

Yeah, that's right.

OpenStudy (anonymous):

sweet!!! thanks!

OpenStudy (anonymous):

please review the steps and make sure you understand what was done, and why.

OpenStudy (anonymous):

I shall try!!

OpenStudy (anonymous):

it seems simple enough, it's just remembering all the different steps for different types of problems that's difficult. :)